Solve each of the following problems.
1. Find the Harmonic mean between 1/2 and 1/10
2. Determine the harmonic mean between x and y

Answer by Theo(13342)

here's a reference.

https://www.mathsisfun.com/numbers/harmonic-mean.html

the basic procedure is:

find the reciprocals of the numbers involved.

find the average of those reciprocals.

take the reciprocal of that.

that's your harmonic mean.

in your problem.

the numbers are 1/2 and 1/20.

their reciprocals are 2 and 20.

the average of those reciprocals is 22/2 = 11

the reciprocal of that is 1/11.

that's your harmonic mean.

it's decimal equivalent is .090909091 when rounded to 9 decimal points.

the harmonic mean of x and y is as shown below.

the reciprocal of x and y is 1/x and 1/y.

the sum of 1/x and 1/y is 1/x + 1/y.

1/x * y/y = y/xy.

1/y * x/x = x/xy.

their sum is (x + y) / xy.

their average is (x + y) / xy * 1/2 = (x + y) / 2xy.

the reciprocal of their average is 2xy / (x + y).

that's your harmonic mean of x and y.

as an example, if x = 3 and y = 5, then their harmonic mean would be equal to:

(2*3*5)/(3+5) = 30/8 = 3.75.

if you looked for their harmonic mean directly, you would:

take the reciprocal of each of those numbers = 1/3 and 1/5.

add them up to get 1/3 + 1/5 = 5/15 + 3/15 = 8/15

divide that by the number of numbers = 2 to get (8/15)/2 = 8/30

take the reciprocal of that to get 30/8 which is equal to 3.75.

the solutions to your questions are:

the harmonic mean of 1/2 and 1/20 is equal to 1/11.

the harmonic mean of x and y is 2xy/(x+y).